OpenStudy (anonymous):

3/48 is what part of 1/12?

OpenStudy (anonymous):

3/48=1/12..

jhonyy9 (jhonyy9):

will be 1 ---- 12 1 48 12*4 4 ------- = ----- * ------ = ------- = ---- 3 12 3 12*3 3 ----- 48 so because 3 4 4 1 ---- * ---- = ---- = ---- 48 3 48 12 hope is understandably